OpenStudy (anonymous):

if sin y= 9/c and tan y= 9/d what is cos y?

jimthompson5910 (jim_thompson5910):

Hint: tan(y) = sin(y)/cos(y)

OpenStudy (anonymous):

@jim_thompson5910 @Hero would it be cos y= d/c

jimthompson5910 (jim_thompson5910):

tan(y) = sin(y)/cos(y) 9/d = (9/c)/cos(y) cos(y)*9/d = 9/c cos(y) = (9/c)/(9/d) cos(y) = (9/c)*(d/9) cos(y) = d/c

jimthompson5910 (jim_thompson5910):

so you got it